OpenStudy (kittiwitti1):

PRE-CALCULUS QUESTIONS. Solved Problems: 1, 2, Since the preview doesn't show equations they are typed below -

OpenStudy (kittiwitti1):

1. Is -4 a zero of \[f(x)=x^5-x^3+2x^2-3x+1\]Show two ways to find the answer. 2. Given a function and one of its zeros, find all the zeros of a function.\[f(x)=x^3-7x^2+17x-15\] 3. State the # of positive/negative real zeros and imaginary zeros for this function\[f(x)=-7x^4+2x^3+x^2-16\]

OpenStudy (anonymous):

One way is to plug it in and see if you get 0.

OpenStudy (anonymous):

another way is to divide by \((x-(-4))\) and see if the result is a polynomial.

OpenStudy (kittiwitti1):

@wio the second could be done by either long or synthetic division, right?

OpenStudy (anonymous):

yeah

OpenStudy (kittiwitti1):

Alright, gotcha. What about #2?

OpenStudy (wolfe8):

What's the zero for number 2?

OpenStudy (kittiwitti1):

Oh, right I forgot to put it up. It's \[2+i\]

OpenStudy (kittiwitti1):

4. Write the function of least degree with integral coefficients that has the following zeros: \[2+/-i\sqrt{3},~-1\] 7. Find all of the zeros of f(x) = 2x^4-9x^3+2x^2+21x-10\]

OpenStudy (wolfe8):

For #2 I think you divide the function with that factor you get: x=2+i x-2-i And solve the the rest equated to 0. Does that make sense?

OpenStudy (kittiwitti1):

I'm trying it with synthetic but I'm stuck...

OpenStudy (wolfe8):

Hmm try: (x-2-i)(x+a)(x+b) expand and use comparison? >.<

OpenStudy (kittiwitti1):

I understand your reasoning but it's not working when I apply it. - Hold on, there are only two minutes left on the computer I'm working at (library) and I have to save my work. I'll come back in about ... an hour

OpenStudy (kittiwitti1):

Back. #7 rewrite: \[f(x)=2x^4-9x^3+2x^2+21x-10\]

OpenStudy (anonymous):

Too tough for me tonight. Good luck.

OpenStudy (kittiwitti1):

LOL. Okay... @douglaswinslowcooper

OpenStudy (lastdaywork):

Q.2 For any polynomial with real coefficients - -complex roots exist in conjugate For any polynomial with rational coefficients - -irrational roots exist in conjugate Now can you solve Q2 ??

OpenStudy (kittiwitti1):

What. @ _ @

OpenStudy (kittiwitti1):

Are you saying that all the roots have "i"?

OpenStudy (lastdaywork):

Nope; what I mean to say is - If (2 + i) is a root of above expression; then (2 - i) would also be one of the roots. (try to guess why? )

OpenStudy (kittiwitti1):

Oh.... right. I forgot the reason... is it because of the complex #?

OpenStudy (lastdaywork):

I am not sure what "complex #" really means :P BTW, this rule will work as long as all the coefficients are real.

OpenStudy (kittiwitti1):

I mean a number with the "i" attached, sorry. Alright... gimme a sec to figure this out lols

OpenStudy (lastdaywork):

Maybe, after this you can teach me #language :D

OpenStudy (kittiwitti1):

XD alrighty

OpenStudy (lastdaywork):

So, can you solve Q2 now ??

OpenStudy (kittiwitti1):

Sorry, hold on

OpenStudy (kittiwitti1):

So the zeros are 2+i and 2-i... and others...?

OpenStudy (kittiwitti1):

I got stuck...

OpenStudy (lastdaywork):

Total number of zeros of a polynomial is equal to the degree of polynomial. As you already have two roots; only 1 remains (and it would be real)

OpenStudy (kittiwitti1):

I can see that but I can't solve :/

OpenStudy (lastdaywork):

Divide f(x) by (x-2-i)(x-2+i) to get a linear equation in f(x). And my previous post was wrong; the third root may be real or may be repeated.

OpenStudy (kittiwitti1):

Okay..... Hold on a sec.

OpenStudy (kittiwitti1):

Almost done. And I booted you up a level on your SmartScore. :p

OpenStudy (lastdaywork):

O yeah.. XD

OpenStudy (kittiwitti1):

Sorry, I'm lagging... hold on ... maybe 20 minutes

OpenStudy (lastdaywork):

Never mind; take you time :)

OpenStudy (kittiwitti1):

(2+i)(2-i)= 4 -2i + 2i - 1 = 3?

OpenStudy (lastdaywork):

Actually - (2+i)(2-i)= 4 -2i + 2i - i^2 = 4 + 1 = 5 In general; you can write - (a + ib)(a - ib) = a^2 + b^2

OpenStudy (kittiwitti1):

o.o oh, right. Hmmm... okay

OpenStudy (kittiwitti1):

Alright, so then I use synthetic division with the 5?

OpenStudy (lastdaywork):

No; You need to divide f(x) by (x-2-i)(x-2+i) or in other words; you need to divide f(x) by (x^2 - 4x + 5)

OpenStudy (kittiwitti1):

Oh, okay.

OpenStudy (kittiwitti1):

... I have no idea how to divide two polynomials.

OpenStudy (kittiwitti1):

Do I simplify into factors and then simplify that?

OpenStudy (lastdaywork):